# On the reconstruction map in JT gravity

### v1

@@ 3 Finding $R^{*}$ for JT gravity
Returning to JT gravity, we have a holographic map $V:\mathcal{H}_{0}\to\mathcal{H}$ given by
$V\ket{x}_{0}=\mathrm{e}^{-S_{0}/2}\sum_{i=0}^{\infty}\phi_{E_{i}}(x)\ket{E_{i}}~{},$ (3.1)
@@
Let us start to make this more precise. By definition, the reconstruction of an operator $O$ on a state |$\ket{\psi}$ has error $\varepsilon$ if
$\lVert R^{*}(O)V\ket{\psi}-VO\ket{\psi}\rVert<\varepsilon~{}.$ (3.2)
Because $V$ is diagonal in the energy basis [1], it follows that the effective theory Hamiltonian $H_{0}$ is perfectly reconstructed by the fundamental theory Hamiltonian $H$ on all states, which encourages us to define $R^{*}$ to satisfy
@@
Unfortunately, this runs into two complications. One [5] is that $\delta$ is technically not a well-defined self-adjoint operator because $H_{0}$ is bounded below, but let us initially proceed as though it were, to illustrate the idea. Another is that in the fundamental theory, $H$ has an irregular random matrix spectrum, and therefore cannot generate perfect shifts of an orthonormal basis of $\delta$-eigenstates. We will have to address these issues by forming an approximate conjugate relationship between $H$ and $R^{*}(\delta)$ that upholds our guiding principle, using tools from quantum ergodicity [4]. The rest of this section will explain how to do this more carefully.
We want to decide what operator in the fundamental theory should be $R^{*}(\delta)$. A priori there are many options, but our guiding principle instructs us to look for an operator that minimizes the reconstruction error on semiclassical states. For definiteness, let us consider |$\ket{\psi}\in\mathcal{H}_{0}$ to be the Hartle-Hawking state projected onto a microcanonical window of width $\Delta E\sim O(1)$, and let |$\ket{\psi(t)}:=\mathrm{e}^{-\mathrm{i}H_{0}t}\ket{\psi}$. We will regard these |$\ket{\psi(t)}$ as our semiclassical states, at least for $t$ small compared to the Heisenberg time. We would like to find an operator $R^{*}(\delta)$ in the fundamental theory that satisfies
$\lVert R^{*}(\delta)V\ket{\psi(t)}-V\delta\ket{\psi(t)}\rVert<\varepsilon$ (3.5)
@@ 3.1 Action-angle variables in the effective theory
As reviewed in the introduction and derived in [5], the phase space of classical JT gravity is $\mathbb{R}^{2}$ with canonical symplectic form
@@
All time-independent classical Hamiltonian systems on a two-dimensional phase space are (locally) integrable. We can then look for “action-angle” variables where the integrability is manifest, and $H$ only depends on the momentum (action) variable. This has been achieved in [5]. Define the canonical variables
@@
Classical trajectories that are incoming from $x=+\infty$ with $p<0$ have $\widetilde{X}\rightarrow-\infty$, while the outgoing trajectory with $p>0$ has $\widetilde{X}\rightarrow+\infty$.
In a system with closed orbits, the action-angle variables $(\widetilde{P},\widetilde{X})$ would be uniquely defined by the periodicity requirement $\widetilde{X}\sim\widetilde{X}+2\pi$. In classical JT gravity, we do not have closed orbits. This means that there is an ambiguity in the action-angle variables. An alternative choice
@@
The effective theory $\mathrm{JT}_{\infty}$ comes from imposing the canonical commutation relations $[\hat{x},\hat{p}]=\mathrm{i}$. We will define action and angle operators $\hat{P}$ and $\hat{X}$ by their classical relationship to $x,p$ and promoting $x,p$ to operators:
$\begin{split}\hat{P}&=P(\hat{x},\hat{p})\\ \hat{X}&=X(\hat{x},\hat{p})\end{split}$ (3.16)
There are operator ordering ambiguities to worry about, but we will argue later that they lead to only small differences in the questions we study. Note here we temporarily wrote hats for clarity, but from now on will go back to not using them.
At the classical level, we stress that there is no reason to prefer one $(X,P)$ action-angle pair over any other. But, when we build a reconstruction map in Section 3.3, we will explain why there is a preferred choice of $F$ in (3.15) that correctly reproduces the spectrum of the fundamental theory.
@@ 3.2 Action-angle operators in the fundamental theory
We will also call $N$ the “number operator” and $\mathrm{e}^{\pm\mathrm{i}\theta}$ the raising/lowering operators. Here |$\ket{-1}=0$ is a null vector. Note this implies
@@
which is similar to the harmonic oscillator algebra, however here $\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{\mathrm{-i}\theta}\neq N$ (where $\mathrm{e}^{\mathrm{i}\theta}=(\mathrm{e}^{\mathrm{-i}\theta})^{\dagger})$. Also note that since $N$ is not directly given in terms of $\mathrm{e}^{\pm\mathrm{i}\theta}$, the formula for the raising and lowering operators are modified from the textbook form for the harmonic oscillator. Note too that $[\mathrm{e}^{\mathrm{i}\theta},\mathrm{e}^{-\mathrm{i}\theta}]\neq 0$ because $\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{-\mathrm{i}\theta}|0\rangle=0$.
We emphasize again that we have not yet related the |$\ket{n}$ basis to the energy eigenbasis |$\ket{E_{n}}$. Different relationships define different pairs of operators $N,\mathrm{e}^{\pm\mathrm{i}\theta}$. We have left this unfixed so that we can determine which choice best upholds our guiding principle. That is, in the next section we will consider reconstructing $P$ and $X$ roughly as $N$ and $\theta$, and then ask which particular $N$ and $\theta$ make the reconstruction error the smallest at early times. We will claim (and argue in Appendix A) that the answer is to line up the bases of $N$ and $H$, identifying |$\ket{n}=\ket{E_{n}}$.
Let us comment on why we have defined $\mathrm{e}^{\pm\mathrm{i}\theta}$ instead of $\theta$ directly. While naively we would have liked to define a conjugate “angle” operator $\theta$ satisfying $\theta\in(-\pi,\pi]$ and $[N,\theta]=-\mathrm{i}$, unfortunately that periodicity and $N\geq 0$ imply such a $\theta$ would not be a well-defined self-adjoint operator on $\mathcal{H}$. Defining $\mathrm{e}^{\pm\mathrm{i}\theta}$ is a suitable workaround which would imply the canonical commutator if $\theta$ were well-defined, and does correctly account for the periodicity [15, 16].
In the absence of a well-defined $\theta$ operator, it is not possible to define eigenstates of $\theta$ in any direct sense. For example, it was noted in [15, 16] that $\cos\theta$ and $\sin\theta$ are well-defined non-commuting operators, when given by the conventional linear combinations of $\mathrm{e}^{\pm\mathrm{i}\theta}$, so they cannot share an eigenbasis of $\theta$-states. So instead, we treat the angle variable with the following strategy. As with the harmonic oscillator, the lowering operator $\mathrm{e}^{\mathrm{-i}\theta}$ admits (non-normalizable) eigenstates $\theta$ with respective eigenvalues $\mathrm{e}^{\mathrm{-i}\theta}$, $\theta\in(-\pi,\pi]$, which form an overcomplete set of coherent states:
@@ 3.3 Defining the reconstruction map
We now propose a reconstruction map, which we argue best upholds our guiding principle. (The full technical argument can be found in Appendix A.) Loosely speaking, the idea is to define $R^{*}(P)=N$ and $R^{*}(\mathrm{e}^{\mathrm{i}X})=\mathrm{e}^{\mathrm{i}\theta}$. This itself is somewhat too quick though, because it does not fully specify how a general function $f(X)$ is to be reconstructed.
We really define $R^{*}$ as follows. Consider a general operator written as a function $A(X,P)$. Define
$\widetilde{A}(k,P)=\frac{1}{2\pi}\int\limits_{-\pi}^{\pi}\mathrm{d}X\,\mathrm{e}^{-\mathrm{i}kX}A(X,P)~{}.$ (3.22)
and then define the reconstruction map
$R^{*}\left(A(X,P)\right)=\sum_{k=-\infty}^{\infty}\frac{1}{2}\left\{\mathrm{e}^{\mathrm{i}k\theta},\widetilde{A}(k,N)\right\}~{}.$ (3.23)
What we have done is to define the reconstruction map via the Fourier transform, truncating the range of $X$ from (,)($\infty$,$\infty$)( - ∞ , ∞ ) to $[-\pi,\pi)$. Note that this reconstruction map is linear, because the Fourier transform is linear. While this map is not an algebra homomorphism, it is approximately one on very high-energy states: see Appendix B.
@@
To finish defining the reconstruction map, we need to specify exactly which $N,\mathrm{e}^{\mathrm{i}\theta}$ to use, i.e. specify how the basis |$\ket{n}$ relates to the energy eigenbasis |$\ket{E_{n}}$. We argue in Appendix A that there is a “best choice”: align the bases |$\ket{n}=\ket{E_{n}}$. We show that with this choice, our guiding principle is best upheld. This brings along a number of nice features, such as $R^{*}(H_{0})=H$ [^5] One could make other choices, e.g. $N^{\prime}=U_{\Delta n}NU_{\Delta n}^{\dagger}$, where $U_{\Delta n}$ is a block diagonal permutation matrix mixing $|n\rangle$ states within a window of size $\Delta n$. At short times, the predictions of this theory would be close to semiclassical predictions, but not as close as with our reconstruction. This permuted theory would have very different dynamics at late times, however.
@@
Figure 2: A cartoon of the modification of phase space suggested our proposed reconstruction map, which truncates the range of $X$ from (,)($\infty$,$\infty$)( - ∞ , ∞ ) to $[-\pi,\pi)$. In $(x,p)$ coordinates, this deletes the red region, which consists of wormholes long compared to $\mathrm{e}^{S(E)}$. The boundary conditions we place at these new walls can be understood as gluing the remaining phase space in the way depicted, making the paths cyclic.
@@
Notice that $\delta$ essentially reduces to a rescaled version of $\theta$ within narrow energy windows as required. The results of [4] (see Appendix A) then imply $\delta$ is as close as possible to an observable whose shifts are generated by $H$ at early times.
@@ 3.4 Explicit reconstruction of the length operator
with $\rho(E)$ given by (1.20). For a random matrix [17]
@@
This gives [^6] As an aside, we remark that this length operator $R^{*}(x)$ does not satisfy the off-diagonal eigenstate thermalization hypothesis [18], conventionally expected for “physical” observables in a theory with a random matrix spectrum. In particular, off-diagonal matrix elements are not suppressed by $e^{-S(E)/2}$ relative to the diagonal elements in the energy eigenbasis, especially near the diagonal. This is partly due to the simple action-angle dynamics of the effective theory, and implies that the expectation value of length does not settle down to a saturation value in special initial states supported on a small number of energy levels (unlike the “larger” initial states of Sec. 4, in which the length does saturate).
@@
Let us comment on how this is related to the length operators in previous work. The authors of [9, 1] made explicit predictions for the matrix elements of the length operator, at least for $n\neq m$:
@@
In the second line, we have approximated that $1\lesssim|m-n|\ll n$, we have further neglected the tiny fluctuations in $E_{n}$ that encode random matrix statistics, and we have lastly invoked (3.32). Comparing to (3.36), we see that there is a quantitative similarity in the scaling of the matrix elements. A technical difference is that (3.36) vanishes between inequivalent energy eigenstates with $m-n$ an even number, while (3.37) does not.
@@ 4.1 Microcanonical state
where |$\ket{n}$ are the energy eigenstates of the fundamental Hilbert space (3.17). We assume that $d\ll N_{0}$, such that $E_{N_{0}+d}\approx E_{N_{0}}$, and this state is supported in a narrow energy window of size
@@
This state appears identical to the states used to develop the theory of quantum ergodicity [4], which we will use below.
@@
We now can generalize [4] to deduce critical features of of the probability distribution of $x$ in the time-evolved microcanonical state
@@
It follows that the wavefunction remains in a microcanonical “coherent” state $\lvert\tilde{\theta}(t)\rangle$ with overwhelming probability for $t\ll t_{{\rm H}}/\Delta$ (for RMT statistics, this time scale is $t_{{\rm H}}/\sqrt{\ln d}$). This means that the wavepacket remains localized to within $\Delta\vartheta\sim 1/d$ of the classical trajectory $\tilde{\theta}(t)$ until this time, but spreads sharply thereafter, to $\Delta\vartheta\sim 1$ by $t\sim t_{{\rm H}}$. The results discussed thus far are a mild extension of earlier work [4].
@@
The quantitative way that the wave function spreads into an incoherent sum over many different wormhole lengths is also very intriguing. In Appendix A.4, we predict that (for $\mathrm{e}^{S}\ll T\ll\exp[\mathrm{e}^{S}]$, at which time we encounter quantum recurrences [19, 20])
@@
This means that on average, the wormhole has a specific length at very late times, but the wave function explores longer wormholes than average for a noticeable amount of time. What actually happens is that for $t\sim t_{\mathrm{H}}$, the wormhole actually overshoots its final length, before bouncing back: see Figure 1. This is a non-trivial prediction of our formalism. We note an intriguing similarity with predictions of [14], which we will discuss further in the concluding section.

### v2

@@ 3 Finding $R^{*}$ for JT gravity
Returning to JT gravity, we have a holographic map $V:\mathcal{H}_{0}\to\mathcal{H}$ given by [Iliesiu:2024cnh]
[…]
@@
[…]
Because $V$ is diagonal in the energy basis [Iliesiu:2024cnh], it follows that the effective theory Hamiltonian $H_{0}$ is perfectly reconstructed by the fundamental theory Hamiltonian $H$ on all states, which encourages us to define $R^{*}$ to satisfy
@@
Unfortunately, this runs into two complications. One [Harlow:2018tqv] is that $\delta$ is technically not a well-defined self-adjoint operator because $H_{0}$ is bounded below, but let us initially proceed as though it were, to illustrate the idea. Another is that in the fundamental theory, $H$ has an irregular random matrix spectrum, and therefore cannot generate perfect shifts of an orthonormal basis of $\delta$-eigenstates. We will have to address these issues by forming an approximate conjugate relationship between $H$ and $R^{*}(\delta)$ that upholds our guiding principle, using tools from quantum ergodicity [dynamicalqergodicity]. The rest of this section will explain how to do this more carefully.
[…]
@@ 3.1 Action-angle variables in the effective theory
As reviewed in the introduction and derived in [Harlow:2018tqv], the phase space of classical JT gravity is $\mathbb{R}^{2}$ with canonical symplectic form
@@
All time-independent classical Hamiltonian systems on a two-dimensional phase space are (locally) integrable. We can then look for “action-angle” variables where the integrability is manifest, and $H$ only depends on the momentum (action) variable. This has been achieved in [Harlow:2018tqv]. Define the canonical variables
@@
Classical trajectories that are incoming from $x=+\infty$ with $p<0$ have $\widetilde{X}\rightarrow-\infty$, while the outgoing trajectory with $p>0$ has $\widetilde{X}\rightarrow$$\infty$.
[…]
@@
The effective theory $\mathrm{JT}_{\infty}$ comes from imposing the canonical commutation relations $\hat{x},\hat{p}]=\mathrm{i}$. We will define action and angle operators $\hat{P}$ and $\hat{X}$ by their classical relationship to $x,p$ and promoting $x,p$ to operators:
[…]
There are operator ordering ambiguities to worry about, but we will argue later that they lead to only small differences in the questions we study. Note here we temporarily wrote hats for clarity, but from now on will go back to not using them; unless explicitly indicated otherwise, $(P,X)$ will henceforth refer to quantum operators acting on the effective Hilbert space $\mathcal{H}_{0}$.
At the level of the effective theory, we stress that there is no reason to prefer one $(X,P)$ action-angle pair over any other. But, when we build a reconstruction map in Section 3.3, we will explain why there is a preferred choice of $F$ in (3.15) that correctly reproduces the spectrum of the fundamental theory.
@@ 3.2 Action-angle operators in the fundamental theory
[…]
@@
which is similar to the harmonic oscillator algebra, however here $\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{\mathrm{-i}\theta}\neq N$ (where $\mathrm{e}^{\mathrm{i}\theta}=(\mathrm{e}^{\mathrm{-i}\theta})^{\dagger})$. Also note that since $N$ is not directly given in terms of $\mathrm{e}^{\pm\mathrm{i}\theta}$, the formula for the raising and lowering operators are modified from the textbook form for the harmonic oscillator. Note too that $\mathrm{e}^{\mathrm{i}\theta},\mathrm{e}^{-\mathrm{i}\theta}]\neq 0$ because $\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{-\mathrm{i}\theta}|0\rangle=0$.
[…]
Let us comment on why we have defined $\mathrm{e}^{\pm\mathrm{i}\theta}$ instead of $\theta$ directly. While naively we would have liked to define a conjugate “angle” operator $\theta$ satisfying $\theta\in(-\pi,\pi]$ and $N,\theta]=-\mathrm{i}$, unfortunately that periodicity and $N\geq 0$ imply such a $\theta$ would not be a well-defined self-adjoint operator on $\mathcal{H}$. Defining $\mathrm{e}^{\pm\mathrm{i}\theta}$ is a suitable workaround which would imply the canonical commutator if $\theta$ were well-defined, and does correctly account for the periodicity [SusskindGlogowerPhaseOperator, CarruthersNietoPhaseAngle].
In the absence of a well-defined $\theta$ operator, it is not possible to define eigenstates of $\theta$ in any direct sense. For example, it was noted in [SusskindGlogowerPhaseOperator, CarruthersNietoPhaseAngle] that $\cos\theta$ and $\sin\theta$ are well-defined non-commuting operators, when given by the conventional linear combinations of $\mathrm{e}^{\pm\mathrm{i}\theta}$, so they cannot share an eigenbasis of $\theta$-states. So instead, we treat the angle variable with the following strategy. As with the harmonic oscillator, the lowering operator $\mathrm{e}^{\mathrm{-i}\theta}$ admits (non-normalizable) eigenstates $\theta$ with respective eigenvalues $\mathrm{e}^{\mathrm{-i}\theta}$, $\theta\in(-\pi,\pi]$, which form an overcomplete set of coherent states:
@@ 3.3 Defining the reconstruction map
We now propose a reconstruction map, which we argue best upholds our guiding principle. (The full technical argument can be found in Appendix A.) Loosely speaking, the idea is to define $R^{*}(P)=N$ and $R^{*}(\mathrm{e}^{\mathrm{i}X})=\mathrm{e}^{\mathrm{i}\theta}$. We emphasize that these are maps between operators: $(P,X)$ act on the effective Hilbert space $\mathcal{H}_{0}$, while $(N,\theta)$ act on the fundamental Hilbert space $\mathcal{H}$. This itself is somewhat too quick though, because it does not fully specify how a general function $f(X)$ is to be reconstructed.
We really define $R^{*}$ as follows. Consider a general effective theory operator written as a function $A(X,P)$. Let ${A}_{\text{cl}}({X}_{\text{cl}},{P}_{\text{cl}})$ be the classical variable corresponding to this operator (here, we introduce subscripts ${\cdot}_{\text{cl}}$ to differentiate classical variables from effective and fundamental operators). Define the classical function ${\widetilde{A}}_{\text{cl}}({k}_{\text{cl}},{P}_{\text{cl}})$ via the Fourier transform
${\widetilde{A}}_{\text{cl}}({k}_{\text{cl}},{P}_{\text{cl}})=\frac{1}{2\pi}\int\limits_{-\pi}^{\pi}\mathrm{d}{X}_{\text{cl}}\,\mathrm{e}^{-\mathrm{i}{k}_{\text{cl}}{X}_{\text{cl}}}{A}_{\text{cl}}({X}_{\text{cl}},{P}_{\text{cl}})~,$ (3.22)
and then, after promoting ${P}_{\text{cl}}$ to the fundamental theory operator $N$ but retaining ${k}_{\text{cl}}$ as a classical variable, define the reconstruction map on quantum operators $A(X,P)$ in the effective theory as
$R^{*}\left(A(X,P)\right)=\sum_{{k}_{\text{cl}}=-\infty}^{\infty}\frac{1}{2}\left\{\mathrm{e}^{\mathrm{i}{k}_{\text{cl}}\theta},\widetilde{A}({k}_{\text{cl}},N)\right\}~.$ (3.23)
Here, $\widetilde{A}({k}_{\text{cl}},N)$ is a well-defined operator in the fundamental theory via the replacement ${P}_{\text{cl}}\to N$ if, e.g., ${\widetilde{A}}_{\text{cl}}({k}_{\text{cl}},{P}_{\text{cl}})$ is defined through a power series in ${P}_{\text{cl}}$, as it only involves the single operator $N$ and therefore no commutation ambiguities. What we have done is to define the reconstruction map via the Fourier transform, truncating the range of $X$ from $(-\infty,\infty)$ to $\pi,\pi)$. Note that this reconstruction map is linear, because the Fourier transform is linear. While this map is not an algebra homomorphism, it is approximately one on very high-energy states: see Appendix B.
@@
[…]
@@
Figure 2: A cartoon of the modification of phase space suggested by our proposed reconstruction map, which truncates the range of $X$ from $(-\infty,\infty)$ to $\pi,\pi)$. In $(x,p)$ coordinates, this deletes the red region, which consists of wormholes long compared to $\mathrm{e}^{S(E)}$. The boundary conditions we place at these new walls can be understood as gluing the remaining phase space in the way depicted, making the paths cyclic.
@@
Notice that $\delta$ essentially reduces to a rescaled version of $\theta$ within narrow energy windows as required. The results of [dynamicalqergodicity] (see Appendix A) then imply $\delta$ is as close as possible to an observable whose shifts are generated by $H$ at early times.
@@ 3.4 Explicit reconstruction of the length operator
with $\rho(E)$ given by (1.20). For a random matrix [zdbai]
@@
This gives [^6] As an aside, we remark that this length operator $R^{*}(x)$ does not satisfy the off-diagonal eigenstate thermalization hypothesis [DAlessio2016], conventionally expected for “physical” observables in a theory with a random matrix spectrum. In particular, off-diagonal matrix elements are not suppressed by $e^{-S(E)/2}$ relative to the diagonal elements in the energy eigenbasis, especially near the diagonal. This is partly due to the simple action-angle dynamics of the effective theory, and implies that the expectation value of length does not settle down to a saturation value in special initial states supported on a small number of energy levels (unlike the “larger” initial states of Sec. 4, in which the length does saturate).
@@
Let us comment on how this is related to the length operators in previous work. The authors of [Iliesiu:2021ari, Iliesiu:2024cnh] made explicit predictions for the matrix elements of the length operator, at least for $n\neq m$:
@@
In the second line, we have approximated that $1\lesssim|m-n|\ll n$, we have further neglected the tiny fluctuations in $E_{n}$ that encode random matrix statistics, and we have lastly invoked (3.32). Comparing to (3.36), we see that there is a quantitative similarity in the scaling of the matrix elements. A technical difference is that (3.36) vanishes between inequivalent energy eigenstates with mnm-n an even number, while (3.37) does not.
@@
We also comment on a conceptual point regarding the relationship between $R^{*}$ and $V$. Our guiding principle ensures that states of the form (3.1) are fairly narrow wavepackets in the eigenbasis of $R^{*}(x)$ for semiclassical values of $x$, and therefore have expectation values and fluctuations consistent with an eigenstate of the latter to leading order. The crucial requirement for this is that $\phi_{E_{n}}(x)$ has a slowly varying phase factor as a function of $n$, which holds for $x\ll\mathrm{e}^{S}$. Such slowly varying phase factors are also seen for time-evolved semiclassical states well before the Heisenberg time scale, which indeed remain fairly localized in $R^{*}(x)$ due to the slow spreading of wavepackets discussed in Sec. 4.1. However, we do not expect or require that $V\lvert x\rangle_{0}$ is precisely an eigenstate of $R^{*}(x)$, only that it is localized to near the $x$-eigenvalue of the latter if $x\ll\mathrm{e}^{S}$. In essence, we are trading off a precise eigenvalue-eigenstate relationship between $R^{*}$ and $V$ in favor of a guiding principle that allows us to carefully compute nonperturbative quantum dynamics in the fundamental theory.
@@ 4.1 Microcanonical state
[…]
@@
This state appears identical to the states used to develop the theory of quantum ergodicity [dynamicalqergodicity], which we will use below.
@@
We now can generalize [dynamicalqergodicity] to deduce critical features of the probability distribution of $x$ in the time-evolved microcanonical state
@@
It follows that the wavefunction remains in a microcanonical “coherent” state $\lvert\tilde{\theta}(t)\rangle$ with overwhelming probability for $t\ll t_{{\rm H}}/\Delta$ (for RMT statistics, this time scale is $t_{{\rm H}}/\sqrt{\ln d}$). This means that the wavepacket remains localized to within $\Delta\vartheta\sim 1/d$ of the classical trajectory $\tilde{\theta}(t)$ until this time, but spreads sharply thereafter, to $\Delta\vartheta\sim 1$ by $t\sim t_{{\rm H}}$. The results discussed thus far are a mild extension of earlier work [dynamicalqergodicity].
@@
The quantitative way that the wave function spreads into an incoherent sum over many different wormhole lengths is also very intriguing. In Appendix A.4, we predict that (for $\mathrm{e}^{S}\ll T\ll\exp[\mathrm{e}^{S}]$, at which time we encounter quantum recurrences [QuantumRecurrences, BrownSusskind2])
@@
This means that on average, the wormhole has a specific length at very late times, but the wave function explores longer wormholes than average for a noticeable amount of time. What actually happens is that for $t\sim t_{\mathrm{H}}$, the wormhole actually overshoots its final length, before bouncing back: see Figure 1. This is a non-trivial prediction of our formalism. We note an intriguing similarity with predictions of [Balasubramanian:2024lqk], which we will discuss further in the concluding section.
Let us also briefly comment on the difference between our approach and a recent work [Bak:2025eul], which proposed to account for the discrete spectrum of JT gravity by simply modifying the exponential potential of (3.8) with a wall at large $x$, whose growth is chosen to match the spectrum. In this approach, the wormhole growth will begin to deviate from its classical value at times $t\sim\mathrm{e}^{S_{0}}$ – rather than $t\sim\mathrm{e}^{S(E)}/\sqrt{S(E)}$ – which is incompatible with our fundamental postulate.